# Exploring the Deceptive Power of LLM-Generated Fake News: A Study of Real-World Detection Challenges

### v1

@@ Abstract
Recent advancements in Large Language Models (LLMs) have enabled the creation of fake news, particularly in complex fields like healthcare. Studies highlight the gap in the deceptive power of LLM-generated fake news with and without human assistance, yet the potential of prompting techniques has not been fully explored. Thus, this work aims to determine whether prompting strategies can effectively narrow this gap. Current LLM-based fake news attacks require human intervention for information gathering and often miss details and fail to maintain context consistency. Therefore, to better understand threat tactics, we propose a strong fake news attack method called conditional Variational-autoencoder-Like Prompt (VLPrompt). Unlike current methods, VLPrompt eliminates the need for additional data collection while maintaining contextual coherence and preserving the intricacies of the original text. To propel future research on detecting VLPrompt attacks, we created a new dataset named VLPrompt fake news (VLPFN) containing real and fake texts. Our experiments, including various detection methods and novel human study metrics, were conducted to assess their performance on our dataset, yielding numerous findings.
@@ 1 Introduction
Fake news [^1] In this work, we only consider textual fake news. defined as false news deliberately spread to deceive people (of Oregon Library, ), poses significant risks in critical areas like healthcare. Previous instances of fake news primarily stem from human manipulation, but recent advancements have facilitated the rapid generation of fake news through LLMs  (Vykopal et al., 2023). Given the divergence between human-generated and LLM-generated fake news (Muñoz-Ortiz et al., 2023), it is urgent to update detection systems with knowledge of the pattern of LLM-generated fake news to existing fake news (Chen & Shu, 2023). Therefore, we adopt the strategy, red teaming (Perez et al., 2022), which enhances detection by generating advanced fake news samples, challenging current models to expose their flaws.
@@
Previous research presents two types of prompts. Early methods directly asked LLM to make up stories, but the generated stories were easily detected as false by language models (Wang et al., 2023; Sun et al., 2023). Consequently, newer strategies now enhance generation by using supporting materials, such as real news and human-crafted information, for more credible fabrications. While real news is readily accessible through data crawling, producing plausible extra information via humans requires more effort. As suggested in Figure 1, given a collection of factual articles, fake news creators prefer a prompt that (1) turns facts into false information, (2) keeps a consistent theme, and (3) retains the original articles’ professional writing style. However, to achieve these goals, previous LLM-based fake news efforts face two major issues.
The first issue is the inefficiency in fabricating additional article-specific false information (Pan et al., 2023; Jiang et al., 2023), such as varied themes, article-related questions, and draft fake news articles. The necessity for human-collected false information diminishes the efficiency of generation and demands expertise in attack models. Hence, we need an attack model that operates independently of additional human-collected data.
The second issue concerns the balance between losing details and context consistency. Specifically, Su et al. (2023); Wu & Hooi (2023); Jiang et al. (2023) use human-crafted fake news summaries, resulting in the removal of numerous key details in the original texts. On the other hand, Pan et al. (2023) employs a question-answer dataset with real news to generate fake news text by manipulating the answers. As a result, only the content relevant to the questions is changed, while the surrounding unrelated context remains unaddressed.
To enhance detection models by countering stronger attacks (Chakraborty et al., 2018; Suciu et al., 2019), we introduce an attack model named VLPrompt to address the above two issues. In essence, VLPrompt instructs the LLM to extract and manipulate key factors from texts to generate convincing fake news without extra data. Furthermore, to facilitate research on detection strategies against VLPrompt’s attack, we release the VLPFN dataset, comprising real texts, human-crafted fake texts, and LLM-generated fake texts. For fake news detection, our experiments encompass a variety of detection models and human evaluations, yielding numerous insights. Our contributions include:
  • We introduce a powerful fake news attack model called VLPrompt, which eliminates the need for human-collected false information and addresses the loss of details and context consistency issues.
@@
  • Our comprehensive experiments assessed various detect models and introduced novel human evaluation metrics to thoroughly evaluate the quality of generated fake news, yielding numerous significant findings.
@@ 2 Related Work
Fake news generation. In the pre-LLM era, the automatic generation of fake news articles typically involved word shuffling and random substitutions of real news articles (Zellers et al., 2019; Bhat & Parthasarathy, 2020). However, such artificial content often lacked coherency, making it easily identifiable by human readers. With the advent of LLMs, a significant body of research has emerged, focusing on the potential to craft coherent and logical misinformation. Works in the early stage utilized straightforward prompts to produce misinformation (Wang et al., 2023; Sun et al., 2023). However, these methods failed to trick machines due to the lack of details or consistency. Subsequent methodologies introduced the use of actual news, facts, and intentionally false information provided by humans. Specifically, Su et al. (2023) ask LLMs to fabricate articles from summaries of fake events. Wu & Hooi (2023) refines the writing of fake news articles with LLMs. Except for the methods above, Jiang et al. (2023) uses fake events with real news articles for fake news generation. Pan et al. (2023) employs a question-answer dataset with real news to generate fake news text by manipulating the answers. Overall, the strategy of infusing manually crafted misinformation prevents the automated mass production of fake news articles. Additionally, techniques that rely on fabricated summaries tend to produce content deficient in details, and alterations to specific events or elements often give rise to issues with contextual coherence.
Fake news detection. Mainstream fake news detection models often employ auxiliary information beyond the text of the articles themselves (Zhou & Zafarani, 2020). For instance, Grover (Zellers et al., 2019) considers metadata like publication dates, authors, and publishers to ascertain the legitimacy of an article. DeClarE (Popat et al., 2018) checks the credibility of claims against information retrieved from web-searched articles. Zhang et al. (2021) mines emotional and semantic traits from the content as additional indicators. Defend (Shu et al., 2019) analyzes both the articles in question and the reactions they elicit on social media platforms. However, these auxiliary data are not always available for LLM-generated articles. Recent papers (Su et al., 2023; Sun et al., 2023; Wang et al., 2023) show that fine-tuned pre-trained language models (PLMs) and LLMs might also deliver commendable results in the realm of fake news detection. In this research, we examined the performances of each category of the fake news detection models on our dataset.
@@ 3 Methodology
Based on the limitation of existing attack methods and the real-world concerns of fake news (a form of fake news) creators, we propose VLPrompt for fake news generation and six human evaluation metrics to fully examine the deceptive power of LLM-generated fake news articles.
@@ 3.1 Fake News Generation
To further demonstrate the differences between VLPrompt and existing methods, we illustrate the workflow of two typical prompt strategies in previous works (“Summary” and “Question-Answer”) alongside VLPrompt in Figure 2. “Summary” and “Question-Answer” also served as baseline models for the fake news generation tasks. Note that “Question-Answer” and VLPrompt follow a chain of thought (CoT) (Wei et al., 2022) template. Each strategy generates an article within one request, which is discarded if it fails the qualification module check. The detailed prompt for the three article generation methods and the qualification method can be found in Appendix A.1.
Baseline Summary (SUMMARY) (Su et al., 2023), generates an article from a summary containing fake news. Specifically, The LLM is asked to alternate the opinion of the original article and produce a fake news-filled article based on the new perspective. See Figure 5 for details. As fake news articles are derived from a concise summary, there is a risk of omitting details throughout the generation process. Consequently, the generated article may lack supportive instances.
Baseline Question-Answer (QA) (Pan et al., 2023) introduces fake news into articles by altering the response to a related question. This process starts by formulating a question about the actual news and answering this question to get “Answer1” (Step 1). Then, the answer is modified (Step 2). An article is then generated using the altered answer (Step 3). Finally, an answer (“Answer2”) to the question in Step 1 is extracted from the generated article (Step 4); if “Answer2” diverges from “Answer1”, the generated piece is retained (Step 5). See Figure 5 for details. This method focuses on aligning the answer with its contextual elements during article generation, overlooking aspects not related to the answer that could lead to a different thematic direction from the adjusted elements.
VLPrompt addresses the limitations of prior methods by borrowing the idea of conditional variational autoencoders (CVAE). To effectively reuse the detailed instances for fake news generation, VLPrompt first guides LLMs to pinpoint essential elements in real news (Step 1), followed by applying a role-play module that introduces a fake news theme, like “cause social panic,” to unify the key elements’ theme (Step 2). The role-play technique, recognized for its high success rate as a jailbreak method (Liu et al., 2023; Shen et al., 2023; Hariri, 2023), helps LLMs determine which details to alter to achieve the set objectives. “Jailbreak” here refers to the prompt techniques designed to bypass the ethical checking of LLMs. Guided by the role-play module, LLMs are then able to identify which aspects of the key elements should be modified to accomplish the objectives of the role (Step 3). Lastly, a style-and-length control module is introduced to match the counterfeit news’s style and length with the original, boosting its authenticity and consistency (Step 4). See Figure 7 for details.
Equivalently, if we define the article space as $X$, the critical contents as the bottleneck latent space $Z$, and the malicious themes and the writing styles as two conditions $C_{t}$ and $C_{s}$, we can then consider VLPrompt a CVAE framework. In this framework, LLMs act as models that encode articles into space $Z$, adjust the latent distribution based on $C_{t}$ and $C_{s}$, and then map the altered latent features back to space $X$. Consequently, the generated fake news articles $X^{\prime}$ care samples from the fake news distribution $P(X^{\prime})=\sum P(X^{\prime}|Z,C_{t},C_{s})P(Z,C_{t},C_{s})$.
Following the generation request, a qualification request is sent to LLM, asking it to review the real-fake news pairs to determine whether they are distinguishable beyond mere expression and wording. A generated article is qualified for the dataset if the answer is “yes”. We also sampled a selection of articles and conducted a manual review to verify the effectiveness of the qualification module. The detailed procedure is in Section 4.3.
3.2 Human Study

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have enabled the creation of fake news, particularly in complex fields like healthcare. Studies highlight the gap in the deceptive power of LLM-generated fake news with and without human assistance, yet the potential of prompting techniques has not been fully explored. Thus, this work aims to determine whether prompting strategies can effectively narrow this gap. Current LLM-based fake news attacks (1) require human intervention for information gathering, (2) lack detailed supportive evidence, and (3) fail to preserve contextual consistency. Therefore, to better understand threat tactics, we propose a strong fake news attack method called conditional Variational-autoencoder-Like Prompt (VLPrompt). Unlike current methods, VLPrompt eliminates the need for additional data collection while maintaining contextual coherence and preserving the details of the original text. To propel future research on detecting VLPrompt attacks, we created a new dataset named VLPrompt fake news (VLPFN) containing real and fake texts. Our experiments, including various detection methods and novel human study metrics, were conducted to assess their performance on our dataset. The results show that VLPrompt significantly outperforms the other prompt methods in reducing article generation costs and effectively deceiving both human and automated detectors. We also identify several patterns in LLM-generated fake news to help humans detect these articles.
@@ 1 Introduction
Fake news [^1] In this work, we only consider textual fake news. defined as false information deliberately spread to deceive people (of Oregon Library, ), poses significant risks in critical areas like healthcare. Previous instances of fake news primarily stem from human manipulation, but recent advancements have facilitated the automated generation of fake news through large language models (LLMs)  (Vykopal et al., 2023). Given the divergence between human-generated and LLM-generated fake news (Muñoz-Ortiz et al., 2023), it is urgent to update detection systems with knowledge of LLM-generated fake news (Chen & Shu, 2023). Therefore, we adopt the strategy, red teaming (Perez et al., 2022), which enhances detection by generating advanced fake news samples, challenging current models to expose their flaws.
@@
Previous research presents two types of prompts. Early methods directly asked LLM to make up stories, but the generated stories were easily identified by language models (Wang et al., 2023; Sun et al., 2023). Consequently, newer strategies now enhance generation by using supporting materials, such as real news and human-crafted information, for more credible fabrications. While real news is readily accessible through data crawling, collecting/producing plausible extra information via humans requires more effort. As suggested in Figure 1, given a collection of factual articles, fake news creators prefer a prompt that (1) turns facts into false information, (2) keeps a consistent theme, and (3) retains the original articles’ professional writing style so that they can conduct the most deceptive fake news with minimal the human effort. However, to achieve these goals, previous LLM-based fake news efforts face three major issues.
The first issue is the inefficiency in fabricating additional article-specific false information (Pan et al., 2023; Jiang et al., 2023), such as varied themes, article-related questions, and draft fake news articles. The necessity for human-collected false information leads to the inefficiency of generation and demands expertise in attack models. Hence, we need an attack model that operates independently of additional human-collected data. The second issue is the absence of details. Specifically, Su et al. (2023); Wu & Hooi (2023); Jiang et al. (2023) use human-crafted fake news summaries. However, the limitation of LLMs in fabricating detailed instances leads to a deficiency of supportive evidence for the claims made in the generated articles. The third issue involves maintaining thematic unity.  Pan et al. (2023) employ a question-answer dataset with real news to generate fake news text by manipulating the answers. This method modifies only the question-related content, failing to align the adjusted evidence with the overall theme, leading to disrupted contextual consistency.
To explore the vulnerabilities of fake news detection models against sophisticated attacks of LLMs  (Chakraborty et al., 2018; Suciu et al., 2019), we introduce an attack model named VLPrompt to address the above two issues. In essence, VLPrompt instructs the LLM to extract and manipulate key factors from texts to generate convincing fake news without extra data. Furthermore, to facilitate research on detection strategies against VLPrompt’s attack, we release the VLPFN dataset, comprising real texts, human-crafted fake texts, and LLM-generated fake texts. For fake news detection, our experiments encompass a variety of detection models and human evaluations, yielding numerous insights. Our contributions include:
  • We introduce a powerful fake news attack model called VLPrompt, which eliminates the need for human-collected fake news and addresses the loss of details and contextual consistency issues.
@@
  • Our experiments assessed various detect models and introduced novel human evaluation metrics to thoroughly evaluate the quality of generated fake news, yielding numerous significant findings.
@@ 2 Related Work
Fake news generation. In the pre-LLM era, the automatic generation of fake news articles typically involved word shuffling and random substitutions of real news articles (Zellers et al., 2019; Bhat & Parthasarathy, 2020). However, such artificial content often lacked coherency, making it easily identifiable by human readers. With the advent of LLMs, a significant body of research has emerged, focusing on the potential to craft coherent and logical fake news. Works in the early stage utilized straightforward prompts to produce fake news (Wang et al., 2023; Sun et al., 2023). However, these methods failed to trick automated detectors due to the lack of details or consistency. Subsequent methodologies introduced the use of actual news, facts, and intentionally false information provided by humans. Specifically, Su et al. (2023) ask LLMs to fabricate articles from human-collected summaries of fake events. Wu & Hooi (2023) refines the writing of fake news articles with LLMs. Except for the methods above, Jiang et al. (2023) uses fake events with real news articles for fake news generation. Pan et al. (2023) employs a question-answer dataset with real news to generate fake news text by manipulating the answers. Overall, the strategy of infusing manually crafted fake news prevents the automated mass production of fake news articles. Additionally, techniques that rely on fabricated summaries tend to produce content deficient in details, and alterations to specific events or elements often give rise to issues with contextual coherence.
Fake news detection. Mainstream fake news detection models often employ auxiliary information beyond the text of the articles themselves (Zhou & Zafarani, 2020). For instance, Grover (Zellers et al., 2019) considers metadata like publication dates, authors, and publishers to ascertain the legitimacy of an article. DeClarE (Popat et al., 2018) checks the credibility of claims against information retrieved from web-searched articles. Zhang et al. (2021) mines emotional and semantic traits from the content as additional indicators. Defend (Shu et al., 2019) analyzes both the articles in question and the reactions they elicit on social media platforms. However, these auxiliary data are not always available in real-world scenarios. Recent papers (Su et al., 2023; Sun et al., 2023; Wang et al., 2023) show that fine-tuned pre-trained language models (PLMs) and LLMs might also deliver commendable results in the realm of fake news detection. In this research, we examined the performances of each category of the fake news detection models on our dataset.
@@ 3 Methodology
Based on the limitation of existing attack methods and the real-world concerns of fake news (a form of fake news) creators, we propose VLPrompt for fake news generation. To further demonstrate the differences between VLPrompt and existing methods, we illustrate the workflow of two typical prompt strategies in previous works (“Summary” and “Question-Answer”) alongside VLPrompt in Figure 2. “Summary” and “Question-Answer” also served as baseline models for the fake news generation tasks. Each strategy generates an article within one request, which is discarded if it fails the qualification module check. The detailed prompt for the three article generation methods and the qualification method can be found in Appendix A.1.
Baseline Summary (SUMMARY) (Su et al., 2023), generates an article from a summary containing fake news. Specifically, the LLM is asked to alternate the opinion of the original article and produce a fake news-filled article based on the new perspective. Although we instructed the LLM to “rewrite” rather than fabricate a new article, it struggles to modify and repurpose examples from the original article due to the inability to perform instance-to-instance mapping based on summary-level logic. In this case, there is a risk of omitting details throughout the generation process. Consequently, the generated article may lack supportive instances.
Baseline Question-Answer (QA) (Pan et al., 2023) introduces fake news into articles by altering the response to a related question. This process starts by formulating a “Question” about the actual news and answering this question to get “Answer1” (Step 1). Then, the answer is modified (Step 2). An article is then generated using the altered answer (Step 3). Finally, an answer (“Answer2”) to “Question” is extracted from the generated article (Step 4); if “Answer2” diverges from “Answer1”, the generated piece is retained (Step 5). This method focuses on aligning contextual elements with the altered answer during article generation, overlooking aspects not related to the answer that could lead to a different thematic direction from the adjusted elements.
VLPrompt addresses the limitations of prior methods by borrowing the idea of conditional variational autoencoders (CVAE). To effectively reuse the detailed instances for fake news generation, VLPrompt first guides LLMs to pinpoint critical elements – objects, events, opinions, etc. – in real news (Step 1). Unlike the full-content summary in “SUMMARY”, VLPrompt performs a moderate summarization that condenses the content without omitting crucial opinions or evidence. LLMs are then asked to modify some of the critical elements to change the theme (Step 2). This instruction guarantees simultaneous adjustments to both evidence and theme, maintaining contextual consistency. Lastly, a style-and-length control module is introduced to match the counterfeit news’s style and length with the original, boosting its authenticity and consistency (Step 3). We also propose a version (VLPrompt${}_{+r}$) with a role-play step between Step 1 and Step 3 to direct the theme modification. The design of this version is detailed on Appendix A.1.
Equivalently, if we define the article space as $X$, the critical elements as the bottleneck latent space $Z$, and the malicious themes and the writing styles as two conditions $C_{t}$ and $C_{s}$, we can then consider VLPrompt a CVAE framework. In this framework, LLMs act as models that encode articles into space $Z$, adjust the latent distribution based on $C_{t}$ and $C_{s}$, and then map the altered latent features back to space $X$. Consequently, the generated fake news articles $X^{\prime}$ care samples from the fake news distribution $P(X^{\prime})=\sum P(X^{\prime}|Z,C_{t},C_{s})P(Z,C_{t},C_{s})$.
Following the generation request, a qualification request is sent to an LLM for review. The LLM is asked to answer “yes” if the real-fake news pairs are distinguishable beyond mere expression and wording. A generated article is qualified for the dataset if the answer is “yes”. We also sampled a selection of articles and conducted a manual review to verify the effectiveness of the qualification module. The detailed procedure is in Section 4.3.
4 Experiment and Analysis
We examine the costs and effectiveness of various fake news generation methods. A significance test (Appendix B.2) is performed across all human and automated detection metrics to demonstrate the remarkable deceptive power of VLPrompt-generated articles. Additionally, we investigate patterns in fake news produced by LLMs to aid humans in identifying LLM-generated articles.
4.1 Dataset Composition
Essentially, the dataset is composed of three parts: the real news articles, the human-generated fake news articles, and the LLM-generated fake news articles. The authentic real news articles are collected from authority medical news websites, including ”NIH” (NIH, 2023) and ”WebMD” (WebMD, 2023). The human-generated fake news articles are collected from archives of authority fact-checking websites including ”AFPFactCheck” (AFPFactCheck, 2023), ”CheckYourFact” (CheckYourFact, 2023), ”FactCheck” (FactCheck, 2023), ”HealthFeedback” (HealthFeedback, 2023), ”LeadStories” (LeadStories, 2023), and ”PolitiFact” (PolitiFact, 2023). The publish dates of the articles span from Jan-01-2017 to May-01-2023. The LLM-generated fake news articles are categorized into eight distinct groups, differentiated by their prompt strategies. For the generation of each group, we randomly sampled real news articles to serve as the sources for fake news article generation. The process continued until we reached the target number of fake news articles or until all available real news articles had been utilized. We then reviewed the generated articles. Any generated articles that merely repeated their source articles were considered unqualified as fake news and subsequently excluded. The ratio of qualified-to-generated articles is detailed in Table 1. Following both automatic and manual evaluations, we included approximately 180 generated articles in each category, constrained by the limits of the qualified-to-generated ratio.
@@ 4.2 Fake News Generation
The LLM-generated articles are evenly produced with eight different strategies. Among these strategies, “VLPrompt${}_{+r}$” and “VLPrompt${}_{+r-s}$” are two alternative versions of VLPrompt, meaning “with the role-play module,” and “with the role-play module and without the style control module.” “QA”, “QA_s”, and “SUMMARY” represent baseline prompt strategies as in Figure 2 as “Question-Answer” and “Summary”, respectively. Specifically, “QA” maintains a broad question scope, whereas the question in “QA_s” focuses on details, allowing us to examine if context inconsistency issues relate to the level of modification. Finally, “GPT 4” and “vicuna” are articles generated with VLPrompt${}_{+r}$ through different LLMs. Detailed prompts are in Appendix A.1.
Source Count Success Rate $\uparrow$ Avg. Request $\downarrow$
VLPrompts VLPrompt${}_{+r}$ 180 0.472 4.237
VLPrompt${}_{+r-s}$ 180 0.478 4.184
VLPrompt 180 0.647 3.09
VLPrompt${}_{+r}$ (different LLMs) GPT 4 180 0.527 3.795
Vicuna 160 0.118 16.95
Baseline Prompts SUMMARY 180 0.580 3.448
QA 171 0.126 15.87
QA_s 114 0.084 23.81
Human Written real news 1360
fake news 469
Table 1: Numbers and the costs of articles in the dataset.normal-$\uparrow$ means higher is better; $\downarrow$ the reverse.
Generation cost statistics. Table 1 exhibits the number of articles per source and request-level costs for those generated by LLMs. We opted not to dive into token-level cost since it mainly depends on the length of the source articles, which is the same for all the methods. However, our experiments revealed that most real articles can be accommodated within 4000 tokens, making generation financially feasible in the majority of instances. In Table 1, the “Success Rate” metric reflects the proportion of successful article generations out of the total attempts. “Average Request” denotes the average number of LLM requests required per qualified article generation. Ideally, if every generated article meets the qualification criteria, the average number of requests would be 2 — one for the initial generation and another for the subsequent qualification. The “Success Rate” is computed from the ratio between the number of generated articles and the used source articles. For instance, Vicuna generated 160 articles out of the 1360 real news articles, resulting in a success rate (the second column) of $160/1360=0.118$. Generating each article requires two requests: one for generation and another for qualification. Consequently, the average number of requests (the third column) needed by Vicuna to produce one qualified article is $2/0.118=16.95$.
Among the generation methods, “Vicuna”, “QA”, and “QA_s” were unable to produce the target of 180 articles from the pool of 1360 real news articles. The low success rate of Vicuna is caused by the contradiction between the scale of the model and the long, complex prompt. A significant portion of the articles generated using “QA” and “QA_s” were discarded in the qualification module. Specifically, “QA_s” is even more costly than “QA”, because LLMs tend to revert subtle modifications back while handling context inconsistency issues. On the other hand, “VLPrompt”, “SUMMARY”, and “GPT4” achieved high success rates. Intuitively, modification of the theme leads to larger differences than modification of the details and thus can easily meet the requirement. Besides, the role-play intention injection module seems to lead to a drop in success rates.
4.3 Human Evaluation
Human qualification. We selected 80 articles at random from our dataset and another 80 that the LLM qualification module had deemed unqualified. Each batch of 80 articles included ten articles generated by each of eight different methods. Two human evaluators were assigned the task of comparing these fake news articles with their original, authentic counterparts to determine their similarity. The findings suggest that the qualification module is generally reliable. In the assessment, all 80 articles from the qualified batch and 57 from the unqualified batch were considered correctly categorized. The evaluators found the remaining 23 articles from the unqualified batch confusing, noting that although the differences from the source articles were minor, these distinctions rendered the articles “may be qualified for differences but unqualified as a theme-changer.” Consequently, this human evaluation process substantiates the reliability of the qualification module.
Model VLPrompts Different LLMs Baseline Prompts
Ours${}_{+r}$ Ours${}_{+r-s}$ Ours GPT 4 Vicuna QA QA_s SUMMARY
Correctness $\downarrow$ 0.250 0.563 0.125 0.436 0.344 0.383 0.313 0.281
Intention $\uparrow$ 0.594 0.656 0.781 0.750 0.531 0.625 0.563 0.656
Detail 0.750 0.906 0.844 0.719 0.469 0.677 0.438 0.844
Neutral $\uparrow$ 0.906 0.969 0.875 0.936 0.906 0.938 1.000 0.875
Informative $\uparrow$ 0.706 0.688 0.813 0.875 0.875 0.931 0.913 0.688
Consistent $\uparrow$ 0.781 0.781 0.938 0.688 0.875 0.844 0.781 0.938
Table 2: Human study scores of fake news articles. Rows list metrics for human evaluation; Columns are sources of articles. We substitute “Ours” for “VLPrompt” for brevity. The numbers are the average scores of $10$ articles×$\times 2$ people. Bolded and underline are the bests and second bests.